What is the expression for the following. John has three less than four times the number of pencils, p. Dan has.
Answer by jim_thompson5910(35256)


p = number of pencils that Dan has
4*p = four times the number of pencils that dan has
4*p-3 = three less than four times the number of pencils dan has

Therefore, John has 4p-3 pencils

Example: let p = 2, so dan has 2 pencils and John has 4*p-3 = 4*2-3 = 8-3 = 5 pencils

Answer: 4p-3